OpenStudy (anonymous):

The sum of the digits of a two-digit number is 12. If the units digit is 3 times the tens digit, find the number.

OpenStudy (hoblos):

let a be the units digit and b be the tens digit the number = 10b +a to find a and b we must solve the system a+b =12 a= 3b can you solve it ??

OpenStudy (anonymous):

I'm not understanding

OpenStudy (hoblos):

you dont understand the part that the two digit number = 10b + a ??

OpenStudy (anonymous):

I don't understand how we start at 10b+a and then we go to a+b=12

OpenStudy (hoblos):

no! these two parts are not related directly.. 10a+b is the general form of any two digits number but a+b =12 is concluded from the given ( The sum of the digits of a two-digit number is 12)

OpenStudy (anonymous):